# Artwork Protection Against Neural Style Transfer Using Locally Adaptive Adversarial Color Attack

### v2

@@ Abstract
Neural style transfer (NST) generates new images by combining the style of one image with the content of another. However, unauthorized NST can exploit artwork, raising concerns about artists’ rights and motivating the development of proactive protection methods. We propose Locally Adaptive Adversarial Color Attack (LAACA), empowering artists to protect their artwork from unauthorized style transfer by processing before public release. By delving into the intricacies of human visual perception and the role of different frequency components, our method strategically introduces frequency-adaptive perturbations in the image. These perturbations significantly degrade the generation quality of NST while maintaining an acceptable level of visual change in the original image, ensuring that potential infringers are discouraged from using the protected artworks, because of its bad NST generation quality. Additionally, existing metrics often overlook the importance of color fidelity in evaluating color-mattered tasks, such as the quality of NST-generated images, which is crucial in the context of artistic works. To comprehensively assess the color-mattered tasks, we propose the Adversarial Color Distance Metric (ACDM), designed to quantify the color difference of images pre- and post-manipulations. Experimental results confirm that attacking NST using LAACA results in visually inferior style transfer, and the ACDM can efficiently measure color-mattered tasks. By providing artists with a tool to safeguard their intellectual property, our work relieves the socio-technical challenges posed by the misuse of NST in the art community.
@@ 1 Introduction
Adversarial attacks [39], a concept primarily explored in machine learning security, have shown promise in subtly altering input data to mislead neural networks. Several studies have demonstrated the effectiveness of adversarial attacks in various domains [1, 12, 24]. Inspired by the disruptive effect of adversarial attacks on machine learning-based systems, we propose to leverage this technique for artwork protection in the context of NST. By strategically embedding specific patterns or "adversarial perturbations" into digital artworks, we aim to systematically disrupt the unauthorized use of original artworks by AI models in advance. This approach offers a more robust and proactive defense mechanism compared to traditional methods like watermarking, as it directly targets the vulnerabilities of neural networks used in NST.
Color plays a crucial role in the perception and aesthetics of visual art [25, 43]. In the context of NST, color consistency is a fundamental aspect of style transfer algorithms [11, 21]. However, most existing reference-based image quality assessment metrics focus on image structure [42] or semantics [15, 35, 44], with limited attention given to color. This oversight leads to a lack of evaluation metrics specifically designed for color-sensitive tasks like NST. To address this issue, we propose the Adversarial Color Distance Metric (ACDM), a novel metric that quantifies the color changes of images after undergoing certain transformations. By capturing color-related properties, ACDM provides a more comprehensive evaluation of color changes between pre- and post-manipulated images, which will also be helpful to exhaustively evaluate the proposed artwork protection method.
@@
  • Through extensive experiments, we demonstrate the effectiveness of LAACA in disrupting NST while maintaining the visual integrity of protected style images, as well as the sensitivity of ACDM for measuring color changes pre- and post-manipulations.
@@ 3 Methodology
In this section, we first define the problem of artwork protection against NST in the adversarial attack framework via a simple yet unified formulation and propose our method. Additionally, we design a color-based metric named Adversarial Color Distance Metric (ACDM) to assess the artistic style difference, which complements the existing Image Quality Assessment (IQA).
@@ 3.1 Problem Definition
Previous unauthorized NST prevention methods, like neural steganography or watermarking, offer post-violation accountability but rely on detecting infringement after the fact. Conversely, the adversarial attack method proactively introduces imperceptible perturbations that degrade NST output quality, deterring potential infringers by rendering the resulting images unsuitable for their intended purpose. the Figure 3 illustrates the difference of those different technique approaches.
Figure 3: The comparison of technique approaches: steganography, watermarker, and adversarial attack.
@@ 3.2 Locally Adaptive Adversarial Color Attack
Inspired by previous works [20, 41], we restrict the perturbation $\boldsymbol{\delta}$ to be frequency-adaptive such that the visual effect is better preserved. This is motivated by the observation that the high correlation between different frequency components and visual effects. This finding guides us to embed adversarial perturbations within the high-frequency areas of the style image. This idea is demonstrated in Figure 4, where restricting perturbations to the high-frequency zone maintains higher visual integrity compared to other patterns. More formally, denote the pixel set of high-frequency components of an image $\boldsymbol{x}$ as $M({\boldsymbol{x}})\subseteq\{(i,j)\mid i=1,\ldots,H,j=1,\ldots,W\}$, and its complement ${{}^{\neg}M(\boldsymbol{x})}$ becomes the corresponding low-frequency pixel set. Thus, the problem is formulated as:
@@
$s.t.\,\|\boldsymbol{\delta}\|_{p}\leq\epsilon\;\text{and }\boldsymbol{\delta}[i,j]=0,\,\text{for }(i,j)\in^{\neg}M(\boldsymbol{x}).$
@@
where $\mathbb{N}$ represents the set of natural number. The pixels in an image $\boldsymbol{x}$ belonging to $M(\boldsymbol{x})$ are denoted as high-frequency zone, and pixels belonging to ${}^{\neg}M(\boldsymbol{x})$ is denoted as low-frequency zone.
@@
Algorithm 1 outlines the protection method transforming $\boldsymbol{x}$ to $\boldsymbol{x}^{\ast}$. We employ an iterative approach with an $\ell_{\infty}$ norm constraint, denoted as $\|\boldsymbol{x}-\boldsymbol{x}^{\ast}\|_{\infty}\leq\epsilon$. before the for loop, we randomly generate a small noise $\boldsymbol{\delta_{0}}$ to avoid the gradient is 0 in loops. For each iteration, the gradient of the loss function with respect to the input is computed, and the perturbations are updated in the direction maximizing the loss value, with a step size of $\alpha$. The perturbations are then clipped to maintain the $\ell_{\infty}$ constraint.
3.3 Adversarial Color Distance Metric
At present, there is no image distance metric specifically designed for distinguishing image differences affected by style from color perspective, i.e., the color consistency is often overlooked in existing research; the commonly used generic metrics are not well-suited for this task either, as evidenced by the experimental results summarized in Table 1. To address this limitation, we propose a new metric, Adversarial Color Distance Metric (ACDM), in the LAB color space, which is constructed to be perceptually uniform and more closely aligns with human visual perception compared to the RGB space [37]. The LAB color space consists of three channels: L represents lightness, ranging from 0 to 100, A and B represent color opponents, with A ranging from -128 ($\bullet$) to 127 ($\bullet$) and B ranging from -128 ($\bullet$) to 127 ($\bullet$).
@@ 3.3 Adversarial Color Distance Metric
$\textnormal{ACDM}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=\sum_{c\in\{L,A,B\}}\mathcal{D}_{c}(\boldsymbol{z}_{1},\boldsymbol{z}_{2}).$ (10)
@@
Consider a toy example of histograms from one channel of 3 images, each with 4 bins. Suppose the normalized histograms of these images are:
$A=[0.0,1.0,0.0,0.0],$
$B=[0.2,0.3,0.5,0.0],$
$C=[0.2,0.3,0.0,0.5].$
Histograms $B$ and $C$ have the same bin values but in a different order. Intuitively, the distribution of $B$ ought to be more similar to $A$ because the position of its 0.5 entry is closer to that of 1.0 in $A$ compared with $C$. However, when we use $\ell_{1}$ loss ($\mathcal{H}_{1}$), $\ell_{2}$ loss ($\mathcal{H}_{2}$) or Cross Entropy ($\mathcal{H}_{3}$) to compare these vectors, we find that $B$ and $C$ have the same loss values when compared to $A$:
$\operatorname{\mathcal{H}_{1}}(A,B)=\operatorname{\mathcal{H}_{1}}(A,C)=0.35,$
$\operatorname{\mathcal{H}_{2}}(A,B)=\operatorname{\mathcal{H}_{2}}(A,C)=0.195,$
$\operatorname{\mathcal{H}_{3}}(A,B)=\operatorname{\mathcal{H}_{3}}(A,C)=1.4437,$
which indicates that although $B$ and $C$ have different similarities to $A$, these three loss functions cannot distinguish between them. On the other hand, if we use EMD to compare these vectors, we find that: $\operatorname{EMD}(A,B)=0.7,\operatorname{EMD}(A,C)=1.2.$
@@ 4.1 Experimental Setup
We use the normalized VGG-19 [11] as our encoder, consistent with NST methods like AdaIN and EFDM, capturing intermediate layer outputs from ${relu1\_1,relu2\_1,relu3\_1,relu4\_1}$. We set $k=4$, $\alpha=8$, $\epsilon=80$, $T=100$ to balance visual integrity and attack effectiveness, with hyperparameter discussions in the ablation studies. Content images are sourced from MS-COCO [23], and style images are from WikiArt [31]. All code are implemented with PyTorch [33], the random seed is fixed as 3407 [34].
@@ 4.4 Ablation Studies
5 Conclusion
In this work, we propose the Locally Adaptive Adversarial Color Attack, a method designed to interrupt unauthorized neural style transfer use cases. Our approach significantly degrades the quality of NST outputs while introducing acceptable perturbations, which will discourage potential infringers from using the protected artwork, because of the bad NST generation. To supplement metrics in evaluating the performance of color-mattered tasks, we introduce the Adversarial Color Distance Metric, which quantifies the color distortions between pre- and post-attack style images. The evaluation of ACDM’s performance in experiments validates its effectiveness in assessing color-related tasks. Experiments demonstrate the efficacy of our attack method in compromising the style transfer process, resulting in significant color distortions and structural dissimilarities in NST images while maintaining the acceptable visual integrity of the post-attack style images. Our approach offers artists a potential tool to protect their intellectual property, with the promise of mitigating or curbing electronic IP infringement. By effectively disrupting unauthorized style transfer, our method is able to serve as a deterrent against the misuse of artists’ stylistic elements without proper attribution or compensation.

### v3

@@ Abstract
Neural style transfer (NST) generates new images by combining the style of one image with the content of another. However, unauthorized NST can exploit artwork, raising concerns about artists’ rights and motivating the development of proactive protection methods. We propose Locally Adaptive Adversarial Color Attack (LAACA), empowering artists to protect their artwork from unauthorized style transfer by processing before public release. By delving into the intricacies of human visual perception and the role of different frequency components, our method strategically introduces frequency-adaptive perturbations in the image. These perturbations significantly degrade the generation quality of NST while maintaining an acceptable level of visual change in the original image, ensuring that potential infringers are discouraged from using the protected artworks, because of its bad NST generation quality. Additionally, existing metrics often overlook the importance of color fidelity in evaluating color-mattered tasks, such as the quality of NST-generated images, which is crucial in the context of artistic works. To comprehensively assess the color-mattered tasks, we propose the Aesthetic Color Distance Metric (ACDM), designed to quantify the color difference of images pre- and post-manipulations. Experimental results confirm that attacking NST using LAACA results in visually inferior style transfer, and the ACDM can efficiently measure color-mattered tasks. By providing artists with a tool to safeguard their intellectual property, our work relieves the socio-technical challenges posed by the misuse of NST in the art community.
@@ 1 Introduction
Adversarial attacks [36], a concept primarily explored in machine learning security, have shown promise in subtly altering input data to mislead neural networks. Several studies have demonstrated the effectiveness of adversarial attacks in various domains [11, 24, 43, 45, 47]. Inspired by the disruptive effect of adversarial attacks on machine learning-based systems, we propose to leverage this technique for artwork protection in the context of NST. By strategically embedding specific patterns or "adversarial perturbations" into digital artworks, we aim to systematically disrupt the unauthorized use of original artworks by AI models in advance. This approach offers a more robust and proactive defense mechanism compared to traditional methods like watermarking, as it directly targets the vulnerabilities of neural networks used in NST.
Color plays a crucial role in the perception and aesthetics of visual art [25, 40]. In the context of NST, color consistency is a fundamental aspect of style transfer algorithms [10, 21]. However, most existing reference-based image quality assessment metrics focus on image structure [39] or semantics [14, 32, 41], with limited attention given to color. This oversight leads to a lack of evaluation metrics specifically designed for color-sensitive tasks like NST. To address this issue, we propose the Aesthetic Color Distance Metric (ACDM), a novel metric that quantifies the color changes of images after undergoing certain transformations. By capturing color-related properties, ACDM provides a more comprehensive evaluation of color changes between pre- and post-manipulated images, which will also be helpful to exhaustively evaluate the proposed artwork protection method.
@@ 3 Methodology
In this section, we first define the problem of artwork protection against NST in the adversarial attack framework via a simple yet unified formulation and propose our method. Additionally, we design a color-based metric named Aesthetic Color Distance Metric (ACDM) to assess the artistic style difference, which complements the existing Image Quality Assessment (IQA).
@@ 3.1 Problem Definition
Previous unauthorized NST prevention methods, like neural steganography or watermarking, offer post-violation accountability but rely on detecting infringement after the fact. Conversely, the adversarial attack method proactively introduces imperceptible perturbations that degrade NST output quality, deterring potential infringers by rendering the resulting images unsuitable for their intended purpose. the supplementary material illustrates the difference of those different technique approaches.
@@ 3.2 Locally Adaptive Adversarial Color Attack
Inspired by previous works [20, 38], we restrict the perturbation $\boldsymbol{\delta}$ to be frequency-adaptive such that the visual effect is better preserved. This is motivated by the observation that the high correlation between different frequency components and visual effects. This finding guides us to embed adversarial perturbations within the high-frequency areas of the style image. This idea is demonstrated in Figure 3, where restricting perturbations to the high-frequency zone maintains higher visual integrity compared to other patterns. More formally, denote the pixel set of high-frequency components of an image $\boldsymbol{x}$ as $M({\boldsymbol{x}})\subseteq\{(i,j)\mid i=1,\ldots,H,j=1,\ldots,W\}$, and its complement ${\neg M(\boldsymbol{x})}$ becomes the corresponding low-frequency pixel set. Thus, the problem is formulated as:
@@
$s.t.\,\|\boldsymbol{\delta}\|_{p}\leq\epsilon\;\text{and }\boldsymbol{\delta}[i,j]=0,\,\text{for }(i,j)\in\neg M(\boldsymbol{x}).$
@@
where $\mathbb{N}$ represents the set of natural number. The pixels in an image $\boldsymbol{x}$ belonging to $M(\boldsymbol{x})$ are denoted as high-frequency zone, and pixels belonging to $\neg M(\boldsymbol{x})$ is denoted as low-frequency zone.
@@
Algorithm 1 outlines the protection method transforming $\boldsymbol{x}$ to $\boldsymbol{x}^{\ast}$. [^1] Our code is available at https://github.com/ZhongliangGuo/LAACA. We employ an iterative approach with an $\ell_{\infty}$ norm constraint, denoted as $\|\boldsymbol{x}-\boldsymbol{x}^{\ast}\|_{\infty}\leq\epsilon$. before the for loop, we randomly generate a small noise $\boldsymbol{\delta_{0}}$ to avoid the gradient is 0 in loops. For each iteration, the gradient of the loss function with respect to the input is computed, and the perturbations are updated in the direction maximizing the loss value, with a step size of $\alpha$. The perturbations are then clipped to maintain the $\ell_{\infty}$ constraint.
3.3 Aesthetic Color Distance Metric
At present, there is no image distance metric specifically designed for distinguishing image differences affected by style from color perspective, i.e., the color consistency is often overlooked in existing research; the commonly used generic metrics are not well-suited for this task either, as evidenced by the experimental results summarized in Table 1. To address this limitation, we propose a new metric, Aesthetic Color Distance Metric (ACDM), [^2] Our code is available at https://github.com/ZhongliangGuo/ACDM. in the LAB color space, which is constructed to be perceptually uniform and more closely aligns with human visual perception compared to the RGB space [34]. The LAB color space consists of three channels: L represents lightness, ranging from 0 to 100, A and B represent color opponents, with A ranging from -128 ($\bullet$) to 127 ($\bullet$) and B ranging from -128 ($\bullet$) to 127 ($\bullet$).
@@ 3.3 Aesthetic Color Distance Metric
$\textnormal{ACDM}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})\quad=\ \sum_{c\in\{L,A,B\}}\mathcal{D}_{c}(\boldsymbol{z}_{1},\boldsymbol{z}_{2}).$ (10)
@@
Consider a toy example of histograms from one channel of 3 images, each with 4 bins. Suppose the normalized histograms of these images are: $A=[0.0,1.0,0.0,0.0]$, $B=[0.2,0.3,0.5,0.0]$, $C=[0.2,0.3,0.0,0.5]$. Histograms $B$ and $C$ have the same bin values but in a different order. Intuitively, the distribution of $B$ ought to be more similar to $A$ because the position of its 0.5 entry is closer to that of 1.0 in $A$ compared with $C$. However, when we use $\ell_{1}$ loss ($\mathcal{H}_{1}$), $\ell_{2}$ loss ($\mathcal{H}_{2}$), Cross Entropy ($\mathcal{H}_{3}$), Cosine Similarity ($\mathcal{H}_{4}$) or Euclidean Distance ($\mathcal{H}_{5}$) to compare these vectors, we find that $B$ and $C$ have the same loss values when compared to $A$: $\operatorname{\mathcal{H}_{1}}(A,B)=\operatorname{\mathcal{H}_{1}}(A,C)=0.35$, $\operatorname{\mathcal{H}_{2}}(A,B)=\operatorname{\mathcal{H}_{2}}(A,C)=0.195$, $\operatorname{\mathcal{H}_{3}}(A,B)=\operatorname{\mathcal{H}_{3}}(A,C)=1.4437$, $\operatorname{\mathcal{H}_{4}}(A,B)=\operatorname{\mathcal{H}_{4}}(A,C)=0.4867$, $\operatorname{\mathcal{H}_{5}}(A,B)=\operatorname{\mathcal{H}_{5}}(A,C)=0.8832$, which indicates that although $B$ and $C$ have different similarities to $A$, these three loss functions cannot distinguish between them. On the other hand, if we use EMD to compare these vectors, we find that: $\operatorname{EMD}(A,B)=0.7,\operatorname{EMD}(A,C)=1.2.$
@@ 4.1 Experimental Setup
We use the normalized VGG-19 [10] as our encoder, consistent with NST methods like AdaIN and EFDM, capturing intermediate layer outputs from ${relu1\_1,relu2\_1,relu3\_1,relu4\_1}$. We set $k=4$, $\alpha=8$, $\epsilon=80$, $T=100$ to balance visual integrity and attack effectiveness, with hyperparameter discussions in the ablation studies. Content images are sourced from MS-COCO [23], and style images are from WikiArt [18].
@@ 4.4 Ablation Studies
5 Conclusion, Limitations, and Future Work
In this work, we propose the Locally Adaptive Adversarial Color Attack, a method designed to interrupt unauthorized neural style transfer use cases. Our approach significantly degrades the quality of NST outputs while introducing acceptable perturbations, which will discourage potential infringers from using the protected artwork, because of the bad NST generation. To supplement metrics in evaluating the performance of color-mattered tasks, we introduce an IQA, ACDM, which quantifies the color distortions between pre- and post-attack style images. The evaluation of ACDM’s performance in experiments validates its effectiveness in assessing color-related tasks. Experiments demonstrate the efficacy of our attack method in compromising the style transfer process, resulting in significant color distortions and structural dissimilarities in NST images while maintaining the acceptable visual integrity of the post-attack style images. However, our work has some limitations. The method’s runtime on common GPUs is not optimal, and the numerous hyperparameters currently provided may not be suitable for all images. Categorizing style images into abstract and realistic paintings, our method is more effective on abstract paintings, possibly because their bolder colors and high-frequency components provide a larger manipulation space. Future work will focus on improving computational efficiency and exploring ways to make these parameters adaptive. We aim to enhance the method’s applicability across a wider range of images and hardware configurations. Overall, our approach offers artists a potential tool to protect their intellectual property, with the promise of mitigating or curbing electronic IP infringement.